OpenStudy (anonymous):

Use substitution to determine if the ordered pair is an element of the solution set of the system of equations. (4,-2) {x+y=2 {y+2x=5

OpenStudy (anonymous):

you can just plug them in?

OpenStudy (anonymous):

that's what I did,but my teacher said I was wrong...

OpenStudy (anonymous):

(A) x+y=2 (B) y+2x=5 x=2-y (from equation A) take this sub into equa B y+2(2-y)=5 y+4-2y=5 -y+4=5 y=-1 x=3

OpenStudy (anonymous):

hence (4,-2) is not a solution since it is not equal to (3,-1) and nor is it a multiple of (3,-1)

OpenStudy (anonymous):

oh okay thank you , can you help me with a similar one please?